3) A 4.33 g sample of calcium sulfide was decomposed into its constituent elements, producing 2.41 g of calcium and 1.92 g of sulfur. Which of the statements are consistent with the law of constant composition (definite proportions)?

a)The mass percentage of calcium plus the mass percentage of sulfur in every sample of calcium sulfide equals 100%.

b)The mass ratio of S to Ca in every sample of calcium sulfide is 0.80.

c)The ratio of calcium to sulfur will vary based on how the sample was prepared.

d)Every sample of calcium sulfide will have 55.6% mass of calcium.

e)Every sample of calcium sulfide will have 2.41 g of calcium.

2)*Use the Conservation of Mass Interactive to answer the question.*

Consider the reaction: MnS(s)+2HCl(aq)⟶MnCl2(aq)+H2S(g)

If the small amount of MnS is combined with the large amount of HCl, what is the mass of the products? The mass of any excess reactants will also be included in this total. mass: ___g
Answer #1

3) Mass of CaS = 40 + 32 = 72

Mass % of Ca = 40 / 72 * 100 = 55.56%

Mass % of S = 32/72 * 100 = 44.44%

a)True. As seen above, the sum of mass perecnt of Ca and S will always be 100%

b) Mass ratio = 32 / 40 = 0.8

True

c) False. The law of constant proportions states that CaS will always have the same ratio of Ca and S by mass.

d) True as proved above

e) False. Only a 4.33 g of CaS will have this amount of Calcium